OpenStudy (anonymous):

The value, V , of a Tiffany lamp, worth $225 in 1965, increases at 15% per year. Its value in dollars t years after 1965 is given by V=225(1.15)^t Find the average value of the lamp over the period 1965-1993 (INTEGRAL PROBLEM)

OpenStudy (amistre64):

225 is a constant, you can pull that aside and focus on 1.15^t

OpenStudy (amistre64):

ln1.15(x) = t try to substitute that

OpenStudy (amistre64):

not ln...but log1.15(x) = t

OpenStudy (anonymous):

why did you take though

OpenStudy (amistre64):

?

OpenStudy (anonymous):

log ...im sorry

OpenStudy (amistre64):

just thought itd be easier to play with.... I got no objection to another way :)

OpenStudy (anonymous):

oh ok...is there a general pattern that i should follow...or even an order

OpenStudy (amistre64):

i think I recall the base other than e stuff now...

OpenStudy (amistre64):

a pattern..... nothing concrete that I am aware of, just find a function that you can work with easily

OpenStudy (amistre64):

1.15^t is the derivative of some F(x) what do we know about exponent derivatives?

OpenStudy (anonymous):

OMG im so confused

OpenStudy (amistre64):

Dx(5^x) = Dx 5^x right? or am i forgetting something there...which I am prone to do

OpenStudy (amistre64):

you want to integrate this function right?

OpenStudy (anonymous):

yea

OpenStudy (amistre64):

then we need to "play" around with it to get it into some form that we can work with or recognize that is easy for us right?

OpenStudy (anonymous):

oh ok

OpenStudy (amistre64):

we dont want to change any value to it, just the way it "looks"

OpenStudy (anonymous):

alright... so its log 1.15(x)= t {where did u get x from}

OpenStudy (amistre64):

from the garbage can inthe back of my memory :) i think we should try it another way... I had forgotten about the derivatives or exponential functions other then base "e"

OpenStudy (anonymous):

lol..ok

OpenStudy (amistre64):

lets try to "derive" a function down to something that looks close to our problem, that might help us "remember" thru the cobwebs :)

OpenStudy (amistre64):

y = 5^x how would we "derive" this? do you recall?

OpenStudy (anonymous):

lnx (5^x)

OpenStudy (amistre64):

you sure? cause it looks right to me...

OpenStudy (anonymous):

haha.. i think ...wait let me check

OpenStudy (amistre64):

remember the steps to get it?

OpenStudy (anonymous):

what steps?

OpenStudy (anonymous):

it actually is ln(5) 5^x

OpenStudy (amistre64):

thats better :)

OpenStudy (anonymous):

so damn close

OpenStudy (amistre64):

ok..now were ready lol

OpenStudy (anonymous):

thanks for not bailing on me half way

OpenStudy (amistre64):

so the derivative of 5^t would be?

OpenStudy (amistre64):

i got nowhere to go ....

OpenStudy (anonymous):

ln(5)5^t

OpenStudy (amistre64):

and that is our "key" to this problem

OpenStudy (amistre64):

if we can get it to "look" like ln(1.15) 1.15^t we can easily integrate it back up right? so how do we change the way something looks without cahnge the value of it?

OpenStudy (amistre64):

14 times what = 14?

OpenStudy (anonymous):

*1

OpenStudy (amistre64):

i knew that number would be useful :) we need a convenient form of the number "1". that has ln(1.15) in it...what would that be?

OpenStudy (anonymous):

omg...idk

OpenStudy (amistre64):

lol ..... what do you know that equals 1?

OpenStudy (amistre64):

does 4/2 = 1? does 4/3 =1? does 4/4 =1??

OpenStudy (anonymous):

ohhh ....4/4?

OpenStudy (amistre64):

instead of a "4" we need what..... ln(1.15) right?

OpenStudy (anonymous):

yea

OpenStudy (amistre64):

good, lets set up out integrand then...

OpenStudy (amistre64):

this is our starting point \[\int\limits_{} 225(1.15^t)\]

OpenStudy (anonymous):

dont we have to use F(b)-F(a)

OpenStudy (amistre64):

lets pull out the constants to get this:\[225 \int\limits_{} 1.15^t\]

OpenStudy (amistre64):

we need to find F(x) before we can use it lol

OpenStudy (anonymous):

dang it

OpenStudy (amistre64):

now lets multiply our integrand by our convenient form of "1"

OpenStudy (amistre64):

\[225 \int\limits_{} [\ln(1.15)/\ln(1.15)] 1.15^t\]

OpenStudy (amistre64):

tell me... is ln(1.15) a constant??

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

then lets pull the bottom one out and leave the top in to use..sound good?

OpenStudy (amistre64):

\[225/\ln(1.15) \int\limits_{} \ln(1.15) 1.15^t\]

OpenStudy (anonymous):

okay that kinda makes sense

OpenStudy (amistre64):

now what do we know our integrand will become?

OpenStudy (anonymous):

oh shoot ......1.15^t

OpenStudy (amistre64):

thats correct :) \[F(t) = [225/\ln(1.15)] 1.15^t \]

OpenStudy (amistre64):

now use your "t" values and figure out the answer :)

OpenStudy (anonymous):

ok...i had one more question....so my values are gonna 1 and 28?

OpenStudy (amistre64):

225 ------- * 1.15^t = F(t) ln(1.15)

OpenStudy (amistre64):

id subtract 1965 from both years to get a span of time

OpenStudy (anonymous):

yeah i subtracted 1965 from 1983

OpenStudy (amistre64):

1983?? or 1993?

OpenStudy (anonymous):

1993...my bad

OpenStudy (amistre64):

it aint 1 to 28, its.... zero to 28. 1965-1965=0

OpenStudy (anonymous):

thats where i was confused .

OpenStudy (amistre64):

1.15^0 = 1 so that aint a biggy

OpenStudy (amistre64):

yeah....it was the subtraction that confused ya :) lol

OpenStudy (anonymous):

OMG...i got it right...THANK YOU

OpenStudy (amistre64):

80599.643 - 1609.88 =?

OpenStudy (amistre64):

yay!!! that means im smart ;)

OpenStudy (amistre64):

we..were smart

OpenStudy (anonymous):

yea...u my friend are the man...

OpenStudy (amistre64):

just dont ask me the hard questions :)

OpenStudy (anonymous):

lol...its all good